# Live Music Models

### v2

@@ C.1 SpectroStream
Here we adopt the recently proposed SpectroStream codec Li et al. (2025), a full-band multi-channel neural audio codec based on residual vector quantization (RVQ) (Zeghidour et al., 2021). Similarly to its predecessor SoundStream Zeghidour et al. (2021), SpectroStream is trained using a combination of adversarial and reconstruction losses. Unlike SoundStream, SpectroStream models audio in the time-frequency domain and adopts a delayed fusion mechanism, which together allow for high-fidelity audio. Specifically, SpectroStream operates on full-bandwidth stereo music at high sample rate ($f_{s}=48$kHz). Relative to other discrete codecs like the Descript Audio Codec Kumar et al. (2023), we train SpectroStream with a relatively slow framerate ($f_{k}=25Hz$) and deeper residual quantizers ($d_{c}=64$), consistent with recommendations from Wu et al. (2025b). Using $10$-bit codebooks ($|\mathbb{V}_{c}|=1024$), this induces an overall bandwidth of 1616kbps. With the goal of facilitating streaming generation via an LM, we reduce the bitrate for generative modeling to 44kbps by generating only the first $16$ RVQ levels (coarse and medium from Figure 3), inducing a throughput target for live generation of $400$ tokens per second.
@@ D.1 Descriptor-based Conditioning
Tempo Beat prediction model Zhao et al. (2022)
@@
We aim to provide precise control over the tempo of the generated stream, defined using beats-per-minute (BPM). Since we do not have access to BPM annotated audio examples, we use an off-the-shelf beat prediction model Zhao et al. (2022) to annotate our dataset, yielding an estimate for beat positions and an averaged BPM value for the entire track. We start by conditioning our model on the target BPM value rounded to the nearest integer using cross-attention.
@@
In addition to style steering using MuLan embeddings, we provide finer grained controls for these features. We use a source separation model Kim et al. (2021) to extract the vocals, bass, drums and other stems from our dataset, and use those stems to generate a set of acoustic features to further condition our model. We specifically extract peak loudness, spectral centroid and bandwidth, chromas and transients separately for every stem.
@@ D.3 Constraining Style Embeddings
Furthermore, biases exist between specific text genres and audio recording quality. To address these issues, impose a latent constraint Engel et al. (2018), by building a dataset comprised of MuLan embeddings from three sources (text, low-quality audio, and high-quality audio) and training a small GAN to transform text based embeddings into high quality audio embeddings using an adversarial setup.
@@ Appendix E Audio Injection
To better control the influence of the user audio on the model output, we use classifier free guidance Ho and Salimans (2021). Specifically, we perform inference on both a positive example where the audio context includes the mixed-in user input, as well as a negative example where the audio context consists of only the model’s output. The final logits used for sampling are calculated as a linear combination of the positive and negative logits, where higher guidance weight ($w$) pushes towards the positive conditioning: $(1+w)\cdot\text{Logits}_{\text{pos}}-w\cdot\text{Logits}_{\text{neg}}$
@@
Appendix F Prompt Transitions

### v3

@@ C.1 SpectroStream
Here we adopt the recently proposed SpectroStream codec [20], a full-band multi-channel neural audio codec based on residual vector quantization (RVQ) [27]. Similarly to its predecessor SoundStream [27], SpectroStream is trained using a combination of adversarial and reconstruction losses. Unlike SoundStream, SpectroStream models audio in the time-frequency domain and adopts a delayed fusion mechanism, which together allow for high-fidelity audio. Specifically, SpectroStream operates on full-bandwidth stereo music at high sample rate ($f_{s}=48$kHz). Relative to other discrete codecs like the Descript Audio Codec [62], we train SpectroStream with a relatively slow framerate ($f_{k}=25Hz$) and deeper residual quantizers ($d_{c}=64$), consistent with recommendations from [63]. Using $10$-bit codebooks ($|\mathbb{V}_{c}|=1024$), this induces an overall bandwidth of 1616kbps. With the goal of facilitating streaming generation via an LM, we reduce the bitrate for generative modeling to 44kbps by generating only the first $16$ RVQ levels (coarse and medium from Figure 3), inducing a throughput target for live generation of $400$ tokens per second.
@@ D.1 Descriptor-based Conditioning
Tempo Beat prediction model [64]
@@
We aim to provide precise control over the tempo of the generated stream, defined using beats-per-minute (BPM). Since we do not have access to BPM annotated audio examples, we use an off-the-shelf beat prediction model [64] to annotate our dataset, yielding an estimate for beat positions and an averaged BPM value for the entire track. We start by conditioning our model on the target BPM value rounded to the nearest integer using cross-attention.
@@
In addition to style steering using MuLan embeddings, we provide finer grained controls for these features. We use a source separation model [65] to extract the vocals, bass, drums and other stems from our dataset, and use those stems to generate a set of acoustic features to further condition our model. We specifically extract peak loudness, spectral centroid and bandwidth, chromas and transients separately for every stem.
@@ D.3 Constraining Style Embeddings
Furthermore, biases exist between specific text genres and audio recording quality. To address these issues, impose a latent constraint [66], by building a dataset comprised of MuLan embeddings from three sources (text, low-quality audio, and high-quality audio) and training a small GAN to transform text based embeddings into high quality audio embeddings using an adversarial setup.
@@ Appendix E Audio Injection
To better control the influence of the user audio on the model output, we use classifier free guidance [38]. Specifically, we perform inference on both a positive example where the audio context includes the mixed-in user input, as well as a negative example where the audio context consists of only the model’s output. The final logits used for sampling are calculated as a linear combination of the positive and negative logits, where higher guidance weight ($w$) pushes towards the positive conditioning: $(1+w)\cdot\text{Logits}_{\text{pos}}-w\cdot\text{Logits}_{\text{neg}}$
@@
Appendix F User Study
To understand how live music models impact users’ creative expression and process, we conducted a study with five music enthusiasts. We were primarily interested in probing how the continually streaming nature of the model impacted their state of flow and creative expression as well as their perceived level of control of the model’s generated music.
Figure 8: Screenshot of the interface used in our study.
User interface
As part of the study, we used the MusicFX DJ interface developed for the Lyria RT model, [^5] http://labs.google/musicfx which includes the advanced controls described in Section D.1. The interface, shown in Figure 8, allows users to input their own natural language prompts (e.g. “oud” or “digital hardcore”), each accompanied by a slider that adjusts its respective influence on the generated music. The system also provides prompt suggestions for inspiration. In addition to steering the model through these prompts, users are also given three high-level controls: density, which adjusts how busy the music should be; brightness, which adjusts the presence of higher frequency sounds; and chaos, which controls how unpredictable the generated music should be. To begin streaming music, users press play, and as they make adjustments to the prompts or high-level controls, the music adjusts accordingly. Finally, users can adjust the key and BPM of the produced music, as well as clear the music history if they want to start fresh.
Study structure
The overall structure of the user study was as follows: (1) participants were given a short demo of the user interface, (2) they then interacted with the interface for 20 minutes, while thinking aloud and explaining their thought process, and (3) afterwards, they were interviewed on their experience. During the task, participants were asked to simply explore the user interface, treat it as an “interactive radio” and discover new sounds (Section F.2). Afterwards, the interview consisted of questions that probed their thoughts on the impact of the model’s continually streaming nature and their perceived level of control (Section F.3). The total time commitment of the study was 30 minutes.
F.1 Findings
Impact of continuous streaming
The continuous nature of the generated music significantly shaped the participants’ creative experience; it gave them something to respond to and improvise with, often drawing comparisons to collaborative and traditional music-making processes. P1 compared interacting with the continual stream to improvising with fellow musicians: “It’s not so different from playing with other people… as you play you react to each other non-verbally to the music. I can see three people playing instruments along with this, steering the model together”. This sentiment was echoed by two other musicians and composers, P4 and P5, who felt the experience mirrored their own composition workflows of creating loops, reacting to them, and incrementally layering additional tracks on software like Ableton. For example, P5 started off with warm acoustic guitar, and then subsequently layered in flute and harp. When asked about her process, P5 explained that the acoustic guitar prompt led to a feeling of “walking in a forest after some rain”, and she added in flute to “mimic the sound of birds chirping”, as well as the harp to further extend the feeling of peace she was building.
A key element of the live model’s continuous stream that users appreciated was its ability to introduce subtle, ongoing variations, even when they had not changed the prompts. This element of gentle evolution was highly valued by P3, who appreciated the opportunity for serendipitous discovery it provided: “I wait to see if the model will produce something interesting… the model might have settled on something that was originally boring, but then jump to something more exciting”. Echoing this point, P4 explained that the subtle variations were “nice in that it [the music] doesn’t feel repetitive”, but at the same time, he noted that these variations would also make him feel like he had “lost something”, if there was something he wanted to keep. Sometimes a melody or texture he liked would suddenly drop out and he felt he had no way to get it back or reinforce it.
On control and influencing the model
In regards to controlling this continuous stream of music, participants felt that the prompts let them broadly steer the model, rather than afford them specific, direct control. P3 felt like he was “guiding and the model is meeting you half way”, while P4 described it as “throwing ingredients in the pot, and the AI is cooking it up”. This perceived level of control, however, changed significantly according to the number of prompts already included. For instance, P1 noted that while the model could be “extremely accurate” with one or two prompts, it became less predictable as more were added. Similarly, P4 felt he had a “good amount of control” when creating an initial sound, but “a lot less control” when trying to make nuanced, real-time adjustments, where additional prompts could suddenly “take over the track” and completely change the sound.
Summary
Overall, the live model’s continuous output fostered a collaborative and improvisational creative process, allowing users to react to and build upon the AI’s output as they would with a fellow musician. While participants appreciated the model’s evolving variations for sparking serendipitous discovery, they ultimately felt more like a guide than a director, possessing the ability to broadly steer the music but lacking the precise control needed to refine it or prevent the loss of desired elements.
F.2 Participant Instructions
We report below the instructions given to participants at the beginning of each session:
Introduction: Hi, thanks for taking the time to participate in this study. During this study, we’ll be creating some sounds with a live, generative music tool. We’ll spend about 15 minutes using the tool, where you’ll be asked to think aloud and explain your thought process as you use it. And afterwards we’ll have a 5-10 minute exit interview on your experience. Any questions so far?
Task: Please point your browser to: https://labs.google/fx/tools/music-fx-dj and share your screen (as well as audio). For the next 20 minutes we’ll be playing around with this tool. As you try out its features, please think-aloud and describe your thought process and reactions. Imagine you’re a composer using this tool to explore new sounds or inspiration for a new project.
F.3 Exit Interview Questions
After completing the task, participants were asked the following questions:
Overall experience: (1) Could you describe your overall experience using the tool? What was surprising, exciting, or frustrating?
Exploration process: (1) Describe a moment when you discovered a particularly interesting or unexpected sound or musical direction. What did you do to get there? (2) What was your thought process behind choosing new prompts or adjusting the sliders? (3) How did the continuous nature of the music influence your exploration or creative process?
Prompt history: (1) How would you describe how the music changed and evolved over the course of the session? (2) How did you feel the history of your prompts shaped the music the model produced later on?
Agency & control: (1) Could you describe the level of control you felt over the music that was generated? (2) How would you describe your partnership with AI in creating the music you heard?
Appendix G Prompt Transitions